OpenStudy (anonymous):

What polynomial has a graph that passes through the given points? (-4,89)(-3,7)(-1,-1)(1,-1)(4,329) how would I start this?

OpenStudy (anonymous):

@amistre64 had a quick and snappy way to do this that i forget you have to set up a system of equations start by calling your polynomial \[p(x)=ax^4+bx^2+cx+d\] and then fill in the numbers you know

OpenStudy (anonymous):

for example since \((1,-1)\) is on the graph, you know that \[p(1)=a+b+c+d=-1\] so there is one equation

OpenStudy (anonymous):

damn typo there i am missing a cube term!

OpenStudy (anonymous):

\[p(x)=ax^4+bx^3+cx^2+dx+e\]\[p(1)=a+b+c+d+e=-1\] \[p(-1)=a-b+c-d+e=1\] etc etc it is a pain

OpenStudy (anonymous):

if it was me, i would cheat and be done in one second

OpenStudy (anonymous):

yeah no kidding.... I

OpenStudy (anonymous):

apparently the answer is \[x^4+2 x^3-3 x^2-2 x+1\]

OpenStudy (mathmale):

Aren't we checking a polynomial equation, not a polynomial expression? Seems to me that with five points on the graph, we'd end up with five equations in five unknowns, which to me signifies y = ax^4 + bx^3 + cx^2 + dx^1 + e (which has five unknown coefficients). So, each of the five points must satisfy the equation y = x4+2x3−3x2−2x+1. Think: how'd you solve a system of five linear equations in five unknowns? Personally, I'd use matrices.